OpenStudy (anonymous):

Find the sum of the first 5 terms of a geometric series whose first term is 3 and whose common ratio is one-fifth. Round your answer to the nearest tenth.

OpenStudy (amistre64):

3, 3r, 3rr, 3rrr, 3rrrr

OpenStudy (amistre64):

geometrics simply multiply the last term by the common ratio to get the next term

OpenStudy (amistre64):

\frac{top}{bottom} if yor trying out the latex coding

OpenStudy (amistre64):

\[\frac{top}{bottom}\]

OpenStudy (anonymous):

amistre there it is given one-fifth

OpenStudy (anonymous):

then 3, 3r,3rr.....hw will dat help

OpenStudy (amistre64):

yes, r = 1/5 the generality of it doesnt change

OpenStudy (anonymous):

ohhh......okie.....so here r=1/5th ?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

thanku...

OpenStudy (amistre64):

your welcome :)